OpenStudy (anonymous):

Find Relative Extrema for: f(x)=1/4 (x+2)(x-1)^2 h(x)=2x^3+5x^2-25x

OpenStudy (math&ing001):

Calculate f'(x) and see when f'(x)=0. If f''(x1)<0 then x1 is a maxima. If f''(x2)>0 then x2 is a minima.

OpenStudy (anonymous):

I'm not even sure what the relative extreme answer should look like

OpenStudy (ranga):

The answer will be an ordered pair such as: (a, b). When x = a, either a minimum or a maximum of f(x) occurs and b is the corresponding minimum or maximum value.

OpenStudy (anonymous):

Ok, but I still don't know how to solve this. Could you help?

OpenStudy (ranga):

f(x)=1/4 (x+2)(x-1)^2 Find the derivative of f(x). f'(x) = ?

OpenStudy (anonymous):

How do I do that

OpenStudy (ranga):

You have not been taught differentiation or finding the derivative yet?

OpenStudy (anonymous):

I probably have, but I must not remember

OpenStudy (ranga):

This may not be the suitable place to teach how to do differentiation. If you have a book you can go through the basics first such as the power rule, the chain rule, the product rule and the quotient rule. Then if you have doubts you can ask and one of us will be able to assist you. But teaching calculus here for the first time via typing will be too time consuming.

OpenStudy (anonymous):

Is the derivative 3x^2-3 / 4

OpenStudy (ranga):

It is 3/4(x - 1)(x + 1) or 3/4x^2 - 3/4.

OpenStudy (anonymous):

Ok so now we know the derivative, how do I find the relative extrema?

OpenStudy (ranga):

Equate it to 0 and solve for x. If you leave the derivative in the factored form: 3/4(x - 1)(x + 1) it will be easier to find the zeros.

OpenStudy (anonymous):

x=1,-1

OpenStudy (ranga):

Yes, those are the x values where a relative minimum or maximum occurs. Put x = 1 and find f(1) Put x = -1 and find f(-1)

OpenStudy (anonymous):

(0,1) is the relative extrema?

OpenStudy (ranga):

The first value should be the x. The second should be the y.

OpenStudy (anonymous):

hmm I'm not sure. When I plugged 1 in I got 0 and when I put -1 I got 1

OpenStudy (ranga):

when x = 1, y = 0. So the point is (1, 0) when x = -1, y = 1. So the other point is (-1, 1) Before concluding these are the extrema (meaning maxima and minima) you should find the second derivative and find evaluate f''(x) at x = 1 and -1. f'(x) = 3/4x^2 - 3/4 f''(x) = ?

OpenStudy (anonymous):

when fx=1, x=1 when fx=-1, x=-1

OpenStudy (ranga):

Not sure what that is.

OpenStudy (ranga):

When x = 1, f(x) = 0 When x = -1, f(x) = 1 which is already done. But find the second derivative of f(x) first. You already have the first derivative as f'(x) = 3/4x^2 - 3/4 Find the derivative again.

OpenStudy (anonymous):

3x/2

OpenStudy (ranga):

Yes. f''(x) = 3x/2 Find f''(x) at x = 1 and -1

OpenStudy (anonymous):

3/2 and -3/2

OpenStudy (anonymous):

i'm sorry I have been doing this study guide for literally 6 hours. These are my last two questions

OpenStudy (ranga):

When x = 1, f''(x) = 3/2. Since f''(3/2) > 0, x = 3/2 is a minimum. When x = -1, f''(x) = -3/2. Since f''(-3/2) < 0, x = -3/2 is a maximum. Now we can say the btwo points we found earlier are the extrema: (1, 0) is a minimum (-1, 1) is a maximum. Since the problem asks for relative extrema you can simply say the extrema are : (-1,1) and (1,0). You can follow the same procedure for the second problem.

OpenStudy (anonymous):

So (-1,1) and (1,0) are the relative extrema?

OpenStudy (ranga):

Yes. Those are the points on the curve f(x) where the function reaches a relative maximum or a minimum.

OpenStudy (anonymous):

For the next one the derivative is 6x^2+10x-25

OpenStudy (ranga):

Yes. Follow the same procedure. Give the final answer and I will check with you later.

OpenStudy (anonymous):

x = 5(-1 sqrt7) / 6, -5(1+sqrt7) / 6

OpenStudy (anonymous):

Thats when x=0

OpenStudy (ranga):

Those are the roots. That is not when x = 0. Those are the x values when h'(x) = 0. It may be better to take the second derivative right away and evaluate h''(x) at those x values. Positive means it is a minimum, negative means it is a maximum. Then you can put the x values in h(x) and find h(x).

OpenStudy (anonymous):

For the local minimum I get 5sqrt7-5 /6 and for the local maximum I get 5sqrt7+5 /6

OpenStudy (anonymous):

Can you please just tell the what the relative extrema is for this last question of my answers are wrong?

OpenStudy (ranga):

The message and reply notifications are not working properly and so unless I visit this page I won't know if you have posted a reply. Let me check your answers...

OpenStudy (ranga):

Doesn't look like your minimum and maximum are correct. Convert the x values where h'(x) = 0 to decimals first and then plug it into h(x).

OpenStudy (anonymous):

26.46 is the maximum decimal

OpenStudy (anonymous):

-26.46 is the minimum decimal

OpenStudy (ranga):

First tell me what your x values are in decimal.

OpenStudy (anonymous):

1.37, -3.04

OpenStudy (anonymous):

I need to have this submitted in the next 15 minutes @ranga

OpenStudy (ranga):

Like I said before, the message and reply notifications are not working properly. Yes those are the correct x values. Put it in h(x) and you should get: 66.02 and -19.72 The extrema are: (1.37, -19.72) and (-3.04, 66.02)